# Physics-Constrained Fine-Tuning of Flow-Matching Models for Generation and Inverse Problems

### v1

@@ 3 Method
In (Domingo-Enrich et al. 2025), Adjoint Matching is introduced as a technique with lower variance and computational cost than standard adjoint methods. The method is based on a Lean Adjoint state, which is initialized as
@@
$\begin{split}\frac{d}{dt}\tilde{a}_{t}&=-\left(\nabla_{\tilde{x}}\tilde{b}^{\text{base}}_{t}(\tilde{X}_{t})^{T}\,\tilde{a}_{t}+\nabla_{\tilde{x}}f(\tilde{X}_{t})^{T}\right)\\ &=-\begin{pmatrix}J_{xx}^{T}&J_{\alpha x}^{T}\\[2.0pt] J_{x\alpha}^{T}&J_{\alpha\alpha}^{T}\end{pmatrix}\begin{pmatrix}a_{t,x}\\[2.0pt] a_{t,\alpha}\end{pmatrix}-\begin{pmatrix}\nabla_{x}f(X_{t},\alpha_{t})^{T}\\[2.0pt] \nabla_{\alpha}f(X_{t},\alpha_{t})^{T}\end{pmatrix}\end{split}$ (2)
where the block-Jacobian is evaluated along the base drift for $X$ and $\alpha$, which means that $J_{ij}=\nabla_{j}\,b^{\text{base}}_{t,i}(X_{t},\alpha_{t})$ for $i,j\in\{x,\alpha\}$. The hyperparameters $\lambda_{x}$ and $\lambda_{\alpha}$ can be used to regulate the extent to which the fine-tuned distribution departs from the base distribution.
The Adjoint Matching objective can then be formulated as a consistency loss:
[…]
$=\tfrac{1}{2}\!\int_{0}^{1}$ $\Bigl{(}\bigl{\lVert}u_{t,x}(X_{t},\alpha_{t})+\sigma(t)\,a_{t,x}\bigr{\rVert}^{2}$
$+$ $\bigl{\lVert}u_{t,\alpha}(X_{t},\alpha_{t})+\sigma(t)\,a_{t,\alpha}\bigr{\rVert}^{2}\Bigr{)}\,dt.$
It can be shown (see (Domingo-Enrich et al. 2025)) that with $f=0$, this objective is consistent with the tilted target distribution for reward r=gr=-g, if optimized with a memoryless noise schedule. This noise schedule is defined as $\sigma(t)=\sqrt{2\eta_{t}}.$ It ensures sufficient mixing during generation such that the final sample $X_{1}$ is independent of $X_{0}$.
Equation 1 is optimized by iteratively sampling trajectories with the fine-tuned model while following a memoryless noise schedule, numerically computing the lean adjoint states by solving the ODE in Equation 2, and taking a gradient descent step to minimize the loss in Equation 3. Note that gradients are only computed through the control $\tilde{u}_{t}$ and not through the adjoint, reducing the optimization target to a simple regression loss. We state the full training algorithm and implementation details in Appendix C.
@@
Figure 3: Parameter-space guidance from sparse observations. The columns on the left show three independent pressure-permeability samples obtained with guidance toward sparse observations on the right. Although the base model was pre-trained solely on observational pressure data, fine-tuning recovers the joint distribution and enables generation of high-quality samples that adhere to the measurements while retaining realistic variability.
@@ 4 Experiments
We demonstrate our method on two PDE systems and cover a broad spectrum of inverse-problem tasks. The experiments span denoising of observational data, guidance from sparse measurements after fine‑tuning, super‑resolution, and the enforcement of reflective and misspecified boundary conditions. Hyperparameters and training specifics are given in Appendix D.
Darcy Flow:
We consider a square domain of $\Omega=[0,1]^{2}$ where a permeability $a(\xi)$ and a forcing term $f(\xi)$ induce a pressure distribution $x(\xi)$. We observe solutions for samples of permeability $a\sim\mu_{a}$, drawn from a discretized Gaussian Process such that $a$ takes values $a\in\{3,12\}$. Pressure $x$ follows the law
$\nabla\cdot\left(a(\xi)\ \nabla x(\xi)\right)-f(\xi)=0,\quad x\in\Omega.$
We usually assume zero Dirichlet boundary conditions. This is a standard dataset used in the Neural Operator literature, established in (Li et al. 2021).
Acoustics:
Here we consider an acoustic wave equation on $\Omega=[0,1]^{2}$ observed for $t\in[0,T]$ with sampled speed of sound $c(\xi)$, representing changing material properties. Again, we sample $c\sim\mu_{c}$ from a discretized 2D Gaussian process with values $c\in\{2,3\}$, which leads to complex acoustic scattering. Boundaries are assumed to be reflective. Starting from an initial condition composed of four isotropic Gaussians equally spaced along both axes, the evolution of pressure is described by:
$\partial_{t}^{2}x(t,\xi)-c(\xi)^{2}\,\Delta x(t,\xi)=0,\quad(\xi,t)\in\Omega\times[0,T].\\$
More detail about the data and its generation is given in Appendix A. For all experiments, we train a FM model on the generated data. As the backbone of the base model we choose a U-FNO (Wen et al. 2022) architecture due to its ability to model sharp edges and local phenomena. Then, before fine-tuning, we pre-train the inverse predictor $\varphi$, mapping final samples $x_{1}$ to parameters $\alpha_{1}$ to minimize the weak residual loss $\mathcal{R}_{\text{weak}}$ on a fixed number of samples generated from the base FM model. As in (Domingo-Enrich et al. 2025), we initialize the fine-tuning model with the same weights as the base model. However, to inform $v_{t,x}^{\text{ft}}$ about the current parameter $\alpha_{t}$, we add conditioning layers to the architecture. Additionally, a separate output head for $v_{t,\alpha}$ is added to the model. During fine-tuning, samples are generated using the memoryless noise schedule. Reported results however are all generated without injected noise, i.e. $\sigma(t)=0$. Details about the implementation can be found in Appendix C.
Observational Data
When working with real-world measured data, we typically face observational noise. In this example, we pre-train a FM model on a noisy version of the Darcy flow dataset, meaning that we add i.i.d noise to each spatial location $\tilde{x}(\xi_{i})=x(\xi_{i})+\sigma\varepsilon_{i}$ with $\varepsilon_{i}\sim\mathcal{N}(0,I)$. During fine-tuning, we minimize the weak PDE residual. This denoising experiment highlights the role of parameter-space regularisation. With a moderate regularisation weight $\lambda_{f}=0.1$ the optimiser suppresses the noise yet stays close to the base permeability, as seen in Fig. 2.
Table 1 compares weak and strong PDE residuals for base FM models trained clean and noisy data respectively, and the corresponding fine-tuned model on the noisy data. For the FM models permeability fields are obtained from the pre-trained inverse predictor $\varphi$. Fine-tuning reduces both residuals substantially demonstrating the efficacy of weak-form optimisation in the presence of noise.
Table 1: Weak- and strong-form PDE residuals for the Darcy denoising experiment averaged over 128 samples. Residuals of the data are given for reference.
Model $\mathcal{R}_{\text{weak}}$ $\mathcal{R}_{\text{strong}}$
Base FM $(4.6\pm 2.0)\times 10^{-1}$ $(8.3\pm 4.2)\times 10^{-1}$
Noisy FM $(9.3\pm 1.5)\times 10^{-1}$ $(1.8\pm 0.2)\times 10^{1}$
Fine-tuned $(2.4\pm 0.6)\times 10^{-1}$ $5.6\pm 0.9$
Noisy data $(9.6\pm 0.7)\times 10^{-1}$ $(2.7\pm 0.3)\times 10^{1}$
Data $(1.0\pm 0.3)\times 10^{-1}$ $1.4\pm 1.0$
To ensure that denoising does not collapse the generative spread we compare mean- and variance-matching scores as well as sample diversity between the base and fine-tuned models (metrics follow (Kerrigan, Migliorini, and Smyth 2023) and are normalised by the data variance, details in Appendix D). With $\lambda_{f}=0.1$ these statistics remain similar to the base model, indicating that the overall distribution is preserved. Without regularisation, diversity in the permeability field drops sharply and the pressure distribution drifts accordingly. Since this result depends on other hyperparameters, such as $\lambda_{x}$, we provide further illustrative results in Appendix E
Table 2: Pointwise statistics indicating sample quality and diversity following (Kerrigan, Migliorini, and Smyth 2023). Based on 256 samples.
Metric Base FM FT FT ($\lambda_{f}=0$)
$\mathrm{MMSE}_{\mathrm{rel}}\,\,x$ ${\times}10^{-2} 1.20×102$${\times}10^{-2}$ $2.42$
$\mathrm{SMSE}_{\mathrm{rel}}\,\,x$ $1.06{\times}10^{-3}$ $1.01{\times}10^{-3}$ $2.71{\times}10^{-2}$
$\mathrm{MMSE}_{\mathrm{rel}}\,\,\alpha$ $2.53{\times}10^{-2}$ $2.35{\times}10^{-2}$ $3.97{\times}10^{-1}$
$\mathrm{SMSE}_{\mathrm{rel}}\,\,\alpha$ $3.13{\times}10^{-2}$ $3.27{\times}10^{-2}$ $1.95{\times}10^{-1}$
$\mathrm{Diversity}_{\mathrm{rel}}\,\,x$ $9.85{\times}10^{-1}$ $9.88{\times}10^{-1}$ $1.21$
$\mathrm{Diversity}_{\mathrm{rel}}\,\,\alpha$ $7.01{\times}10^{-1}$ $6.97{\times}10^{-1}$ $4.29{\times}10^{-1}$
Guidance and Conditional Generation
@@
In many realistic settings dense observations of a state variable are available for pre-training a generative model, whereas only a few measurements of the latent parameter can be collected. To sample from the posterior of parameter–state pairs that respect such sparse evidence we steer the generative process through guidance. Guiding sampling on learned quantities is a common practise for tasks like image inpainting, see (Bansal et al. 2023; Song et al. 2021). For physical data, (Huang et al. 2024) demonstrates guided sampling towards sparse observations from a model that was pre-trained on the joint parameter-state distribution. Our approach applies the same guidance mechanism, however, to a model that was pre-trained on state observations alone.
Let $\alpha^{\star}=\{(\xi_{i},\alpha_{i}^{\star})\}_{i=1}^{m}$ denote the observed parameter values and locations. The normalised mean-squared discrepancy
$\mathcal{L}_{\text{guide}}=\frac{1}{m}\sum_{i=1}^{m}\lvert\hat{\alpha}_{1}^{\text{ft}}(\xi_{i})-\alpha^{*}_{i}\rvert^{2}$
provides a steering signal, which we inject into each Euler step as
$\tilde{x}_{t+\Delta t}=\tilde{x}_{t}+\Delta t\;\tilde{v}_{t}^{\text{ft}}(\tilde{x}_{t})-\zeta\,\Delta t\;\nabla_{\tilde{x}}\mathcal{L}_{\text{guide}}$
with $\zeta>0$ controlling the conditioning strength. Figure 3 confirms that the guided sampler adheres to the sparse measurements while preserving realistic variability. Additional examples are given in Appendix E.
Super-Resolution and Reflective Boundary in 3D
Spatio-temporal acoustic wave propagation is numerically demanding: sharp material discontinuities combined with reflective boundaries produce intricate scattering patterns that are costly to resolve. We represent every space–time pressure field as a three-dimensional sample $x\in\mathbb{R}^{64\times 64\times 64}$ obtained from a low-fidelity solver and train a base Flow-Matching model on this data. At test time the model outputs of pressure and the recovered parameter field $c$ are upsampled by a factor of four through trilinear interpolation and weak PDE residuals and reflective-boundary penalties are evaluated on the $256^{3}$ grid. Because naïve interpolation fails to enforce the zero-normal-gradient boundary condition, fine-tuning is essential. Table 3 shows that fine-tuning is able to significantly reduce the residual of the reflective boundary $\mathcal{R}_{\text{BC}}$, defined as the mean of the squared magnitude of the normal gradient at the boundary, while still adhering to the physical laws imposed by the PDE, indicated by slightly lower weak residuals. We visually present results in Appendix E.
Table 3: Acoustic-wave test case: weak residual $\mathcal{R}_{\text{weak}}$ and reflective-boundary residual $\mathcal{R}_{\text{BC}}$ across 20 generated samples. Fine-tuning sharply reduces the boundary error while further lowering the weak residual.
@@
Model $\mathcal{R}_{\text{weak}}$ $\mathcal{R}_{\text{BC}}$
Base $(4.0\pm 0.8)\times 10^{-3}$ $(2.6\pm 0.1)\times 10^{-1}$
Fine-tuned $\mathbf{(2.2\pm 0.6)\times 10^{-3}}$ $\mathbf{(8.0\pm 0.4)\times 10^{-2}}$
Misspecification
We finally assess the robustness of our fine-tuning procedure to misspecified training data. Starting from the 2-D stationary Darcy setup, we prescribe a non-zero Dirichlet condition on the top $x(\xi_{1},1)=\sin(\pi\,\xi_{1})$, while keeping zero boundary values elsewhere. During fine-tuning, however, we deliberately enforce homogeneous Dirichlet data together with the weak PDE residual. Fig. 4 shows fine-tuning results for different strengths of $\lambda_{x}$ for the same noise sample. As $\lambda_{x}$ increases, the samples conform progressively better to the desired zero boundary without sacrificing physical plausibility, even though the model must extrapolate far beyond its pre-training domain. Owing to parameter regularisation $\lambda_{f}>0$, the fine-tuned solutions still preserve sample-specific permeability features. Additional visualisations, for example with $\lambda_{f}=0$, and a comparison with ECI (Cheng et al. 2024) are provided in Appendix E.
Figure 4: Misspecified-boundary experiment. Each column shows a pressure field generated from the same initial noise but from models fine-tuned with increasing $\lambda_{x}$.

### v2

@@ 3.3 Adjoint Matching
In Domingo-Enrich et al. (2025), Adjoint Matching is introduced as a technique with lower variance and computational cost than standard adjoint methods. The method is based on a Lean Adjoint state, which is initialized as
@@
$\frac{d}{dt}\tilde{a}_{t}=-\left(\nabla_{\tilde{x}}\tilde{b}^{\text{base}}_{t}(\tilde{X}_{t})^{T}\,\tilde{a}_{t}+\nabla_{\tilde{x}}f(\tilde{X}_{t})^{T}\right)=-\begin{pmatrix}J_{xx}^{T}\!&\!J_{\alpha x}^{T}\\[2.0pt] J_{x\alpha}^{T}\!&\!J_{\alpha\alpha}^{T}\end{pmatrix}\begin{pmatrix}a_{t,x}\\[2.0pt] a_{t,\alpha}\end{pmatrix}-\begin{pmatrix}\nabla_{x}f(X_{t},\alpha_{t})^{T}\\[2.0pt] \nabla_{\alpha}f(X_{t},\alpha_{t})^{T}\end{pmatrix}$ (3)
where the block-Jacobian is evaluated along the base drift for $X$ and $\alpha$, which means that $J_{ij}=\nabla_{j}\,b^{\text{base}}_{t,i}(X_{t},\alpha_{t})$} for $i,j\in\{x,\alpha\}$. The hyperparameters $\lambda_{x}$ and $\lambda_{\alpha}$ can be used to regulate the extent to which the fine-tuned distribution departs from the base distribution. The Adjoint Matching objective can then be formulated as a consistency loss:
$\begin{split}\mathcal{L}(\tilde{u};\tilde{X})&=\tfrac{1}{2}\!\int_{0}^{1}\bigl\lVert\tilde{u}_{t}(\tilde{X}_{t})+\sigma(t)\,\tilde{a}_{t}\bigr\rVert^{2}\,dt\\ &=\tfrac{1}{2}\!\int_{0}^{1}\Bigl(\bigl\lVert u_{t,x}(X_{t},\alpha_{t})+\sigma(t)\,a_{t,x}\bigr\rVert^{2}+\bigl\lVert u_{t,\alpha}(X_{t},\alpha_{t})+\sigma(t)\,a_{t,\alpha}\bigr\rVert^{2}\Bigr)\,dt.\end{split}$ (4)
It can be shown (Domingo-Enrich et al., 2025) that with $f=0$, this objective is consistent with the tilted target distribution for reward r=gr=-g, if optimized with a memoryless noise schedule. This schedule ensures sufficient mixing during generation such that the final sample $X_{1}$ is independent of $X_{0}$. To stabilise fine-tuning we introduce a scaled variant of the memoryless noise schedule. Instead of using the canonical choice $\sigma^{2}(t)=2\eta_{t}$ identified by Domingo-Enrich et al. (2025), we adopt
$\sigma^{2}(t)=(1-\kappa)\,2\eta_{t},\qquad 0\leq\kappa<1,$
which retains the theoretical memoryless property (see Lemma 1 in Appendix D.4) while attenuating the magnitude of the noise variance. The introduction of the scaling factor $0\leq\kappa<1$ constitutes a simple but novel extension of the adjoint-matching framework. Whereas prior work highlighted a unique schedule, our analysis shows that a family of scaled schedules remains consistent with the memoryless condition. This additional degree of freedom acts as a numerical stabilisation knob, mitigating blow-ups near $t\to 0$ without losing theoretical consistency. Further, it offers a control–fidelity trade-off by regulating the amount of exploration.
In practice, this flexibility allows practitioners to adapt fine-tuning to the conditioning of the PDE residuals and the stability of the solver, a feature not available in the original formulation.
Equation 2 is optimized by iteratively sampling trajectories with the fine-tuned model while following a memoryless noise schedule, numerically computing the lean adjoint states by solving the ODE in Equation 3, and taking a gradient descent step to minimize the loss in Equation 4. Note that gradients are only computed through the control $\tilde{u}_{t}$ and not through the adjoint, reducing the optimization target to a simple regression loss. We state the full training algorithm and implementation details in Appendix D.5.
@@ 4 Experiments
We evaluate across five settings: four PDE systems (including boundary and system misspecification, and observational noise) and a natural-image model. Unlike latent-space fine-tuning for images, our PDE models operate directly in pixel space. High-variance noise during sampling can drive off-manifold trajectories and perturb PDE residuals, motivating $\kappa>0$ for these models. For base Flow Matching backbones we use U-FNO (Wen et al., 2022) for PDEs and the DiT-based latent FM of Dao et al. (2023) for images. In all experiments we first sample from the base generator and pre-train the inverse predictor $\varphi$ to recover $\alpha$ by minimizing the (PDE) residual, then fine-tune. Following Domingo-Enrich et al. (2025), fine-tuning is initialized from the base weights. We augment capacity to condition $v_{t,x}^{\text{ft}}$ on $\alpha_{t}$ and add a separate head for $v_{t,\alpha}^{\text{ft}}$. Fine-tuning uses a memoryless noise schedule, while all reported results are generated without injected noise ($\sigma(t)=0$). Implementation details appear in App. D.2.
Comparisons, ablations, and metrics.
[…]
4.1 Darcy Flow
Consider a square domain $\Omega=[0,1]^{2}$ where a permeability $\alpha(\xi)$ and forcing $f(\xi)$ induce a pressure field $x(\xi)$ governed by $\nabla\cdot\left(\alpha(\xi)\nabla x(\xi)\right)-f(\xi)=0$ with zero Dirichlet boundary conditions and constant $f$. We draw $\alpha$ from a discretized Gaussian process and corrupt pressures with observation noise before training the base FM. Dataset details are in App. B.
Figure 2 compares three Darcy samples generated from the same noise seed $x_{0}$: the base draw, fine-tuning with our regularization (here $\lambda_{f}=1.0$), and fine-tuning without regularization. The base pressure $x^{\text{base}}$ is visibly contaminated by high-frequency noise, and the inverse predictor $\varphi$ correspondingly yields a scattered, artefact-ridden permeability map $\alpha^{\text{base}}$. With regularization enabled, fine-tuning attenuates noise in the pressure $x^{\text{ft}}$ while remaining close to $\alpha^{\text{base}}$. Because $\alpha^{\text{base}}$ is itself fragmented, some artefacts persist. In contrast, disabling regularization produces a fully denoised pressure and a markedly more coherent $\alpha^{\text{ft}}$, but at the expected expense of erasing sample-specific details present in the base realization.
Figure 2: Darcy denoising (qualitative). Base vs. fine-tuned outputs for a fixed seed. Regularization ($\lambda_{f}=1.0$) denoises while staying close to the base sample. Removing it denoises more aggressively at the cost of fidelity to the base realization. Additional non-curated samples in App. F.3.1. Color maps throughout this work taken from Crameri et al. (2020).
We quantify the controllable trade-offs in Fig. 3. Panel (a) increases $\lambda_{x}=\lambda_{\alpha}$ at $\lambda_{f}=0$, which reduces the PDE residual while also reducing diversity in the inferred permeabilities (measured via the complement of the mean pairwise SSIM; see App. E.1). Panel (b) fixes $\lambda_{x}=\lambda_{\alpha}=20\mathrm{K}$ and varies $\lambda_{f}$, reporting MMDx between the fine-tuned samples and the base dataset. As expected, stronger regularization preserves distributional fidelity (lower MMD) but yields higher residuals. These ablations illustrate how practitioners can target residual reduction or distributional fidelity by tuning $(\lambda_{x},\lambda_{\alpha},\lambda_{f})$.
(a) $\lambda_{x}$ and $\lambda_{\alpha}$ vs. residual/diversity
(b) Regularization $\lambda_{f}$ vs. residual/MMD
Figure 3: Darcy ablations. (a) Increasing $\lambda_{x}=\lambda_{\alpha}$ with $\lambda_{f}=0$ lowers the PDE residual but reduces diversity in the inferred parameters (reported via SSIM-based diversity). (b) Sweeping $\lambda_{f}$ trades PDE residuals against fidelity to the base distribution (MMDx). Each point averages 256 samples with shared noise seeds across settings.
Computationally, adaptation is lightweight: fine-tuning on noisy Darcy requires only 20 gradient steps (hyperparameters in App. E.3) and completes in under 15 minutes on a single NVIDIA L40S, after which sampling proceeds at base-model cost with no inference-time adjustments.
4.2 Guidance on Sparse Observations
@@
In many realistic settings dense observations of a state variable are available for pre-training a generative model, whereas only a few measurements of the latent parameter can be collected. To sample from the posterior of parameter–state pairs that respect such sparse evidence we steer the generative process through guidance. Huang et al. (2024) demonstrate guided sampling towards sparse observations from a model that was pre-trained on the joint parameter-state distribution. Our approach applies a similar guidance mechanism, however, to a model that was pre-trained on noisy state observations alone. We state details on the guiding mechanism in E.4. Figure 4 shows that the guided sampler adheres to sparse measurements while preserving realistic variability in the generated samples. Additional results for different amounts of conditioning observations are given in Appendix F.3.5.
Figure 4: Three samples through guidance towards sparse observations (white markers in right panel) of the permeability, showing a plausible conditional distribution.
4.3 Linear Elasticity
We consider plane–strain linear elasticity on $\Omega=[0,1]^{2}$ with spatially varying Young’s modulus $\alpha(\xi)$ and fixed Poisson ratio. Boundaries are Dirichlet: left/right clamped, top/bottom receive inward sinusoidal normal displacements with zero tangential slip. During fine-tuning, we impose a modified lower-boundary amplitude to induce controlled misspecification (see App. B) and include an MSE boundary penalty in the weak residual. We report quantitative BC results as the MSE at the boundary in Table 1 and a qualitative comparison is provided in App. F. Our method attains low weak/strong residuals while keeping distributional shift modest; PBFM and FM+ECI drift distributionally or present high residuals (full details and non-curated samples in App. E.5, F.3.2).
@@ 4.3 Linear Elasticity
Model BC error (MSE) $\downarrow$ $\mathcal{R}_{\text{weak}}$ (rel) $\downarrow$ $\mathcal{R}_{\text{strong}}$ (rel) $\downarrow$ MMDx $\downarrow$ MMDα $\downarrow$
FM $6.98\times 10^{-5}$ ($\pm$ 0.53) $1.59\times 10^{1}$ ($\pm$ 0.37) $1.83\times 10^{1}$ ($\pm$ 0.66) $0.24$ $0.05$
PBFM $2.32\times 10^{-5}$ ($\pm$ 0.87) $6.32\times 10^{0}$ ($\pm$ 0.82) $4.22\times 10^{0}$ ($\pm$ 0.26) $0.92$ $0.54$
FM+ECI $0.0$ $1.01\times 10^{3}$ ($\pm$ 0.13) $2.49\times 10^{2}$ ($\pm$ 0.32) $1.16$ $0.36$
Ours $1.71\times 10^{-6}$ ($\pm$ 0.50) $6.15\times 10^{0}$ ($\pm$ 0.77) $3.79\times 10^{0}$ ($\pm$ 0.87) $0.15$ $0.12$
Table 1: Linear elasticity under BC misspecification. Quantitative boundary-condition (BC) error, relative weak/strong residuals, and distributional metrics. Our method achieves low residuals with limited distributional shift.
4.4 Helmholtz
We consider time–harmonic wave propagation governed by the heterogeneous Helmholtz equation $\Delta u-(1-i\tan\delta)\,\kappa(x)^{2}u=s$} on $\Omega=[0,1]^{2}$ with Robin boundary conditions. Training data use a small damping term ($\tan\delta>0$) producing complex attenuated fields, while fine-tuning assumes the idealized lossless model ($\tan\delta=0$), inducing a controlled model mismatch.
Model Criterion $\mathcal{R}_{\text{weak}}$ (rel) $\downarrow$ $\mathcal{R}_{\text{strong}}$ (rel) $\downarrow$ MMDx $\downarrow$ MMDα $\downarrow$
FM $1.5\times 10^{1}$ ($\pm$ 0.59) $2.55\times 10^{1}$ ($\pm$ 0.55) $0.18$ $0.03$
PBFM $8.33\times 10^{0}$ ($\pm$ 3.04) $1.22\times 10^{1}$ ($\pm$ 0.33) $0.09$ $0.03$
Base AM $\mathcal{R}_{\text{weak}}$ $4.9\times 10^{0}$ ($\pm$ 1.85) $1.34\times 10^{1}$ ($\pm$ 0.32) $0.15$ $0.04$
Base AM $\mathrm{MMD}_{x}$ $5.64\times 10^{0}$ ($\pm$ 2.09) $1.59\times 10^{1}$ ($\pm$ 0.33) $0.13$ $0.04$
Base AM + $\varphi$ $\mathcal{R}_{\text{weak}}$ $4.99\times 10^{0}$ ($\pm$ 2.12) $1.16\times 10^{1}$ ($\pm$ 0.33) $0.13$ $0.05$
Base AM + $\varphi$ $\mathrm{MMD}_{x}$ $5.46\times 10^{0}$ ($\pm$ 1.94) $1.59\times 10^{1}$ ($\pm$ 0.33) $0.12$ $0.04$
AM $\mathcal{R}_{\text{weak}}$ $4.3\times 10^{0}$ ($\pm$ 1.29) $1.14\times 10^{1}$ ($\pm$ 0.29) $0.07$ $0.04$
AM $\mathrm{MMD}_{x}$ $4.32\times 10^{0}$ ($\pm$ 1.43) $1.05\times 10^{1}$ ($\pm$ 0.30) $0.06$ $0.04$
Table 2: Helmholtz: residuals and distribution (representative configs). Normalized weak/strong residuals and MMD metrics. We include AM variants and a PBFM-style baseline.
[…]
4.5 Stokes Lid-Driven Cavity
We consider steady incompressible flow in the Stokes regime (linear, low–Reynolds-number proxy) governed by $\nabla\!\cdot(\nu(x)\nabla u)+\nabla p=f,\quad\nabla\!\cdot u=0$} with no-slip walls and a smooth moving lid. The dataset uses nonzero Kolmogorov forcing $f\neq 0$, while fine-tuning assumes $f\equiv 0$, creating a systematic model mismatch.
[…]
4.6 Natural Images: Parametric Color Transformation
To demonstrate cross-domain utility, we apply our method to natural images by introducing a parametric recoloring pathway: analogous to the hidden PDE parameter, $\alpha$ here specifies a polynomial color transform that operates outside the latent space, enabling exploration of image appearances not well supported by the base distribution. We use a class-conditional Latent Flow Matching (LFM) model (Dao et al., 2023) pretrained on ImageNet-1k (Deng et al., 2009) and optimize PickScore (Kirstain et al., 2023) with a globally fixed prompt. As a concrete example, we fine-tune on the class macaw with the prompt “close-up Pop Art of a macaw parrot,” yielding the samples in Fig. 6. Joint fine-tuning with recoloring produces markedly more vibrant palettes and, crucially, joint adjustments (e.g., background textures that the recoloring exploits). Details about the recoloring parametrization are given in Appendix E.7 and further non-curated samples are provided in Appendix F.3.6.
(a) Vanilla Adjoint Matching
(b) Joint model with parametric recoloring
Figure 6: Fine-tuning of LFM model on macaw class using prompt “close-up Pop Art of a macaw parrot”, comparing vanilla Adjoint Matching with our joint approach.